1. In Blood Donation Camp, we took a sample of 148 donors, 41 had type A+ blood, 19 had type O+ blood, 28 had type B- blood, 46 had type A-, 10 had type AB+ and 4 had type AB- blood. Set up a frequency distribution and find the following probabilities.

a. A person has type A- blood.
b. A person has type AB+ or type AB- blood.
c. A person has neither type B+ nor type A+ blood.
1
Expert's answer
2021-09-07T23:21:55-0400

Frequency distribution:

Type Number

A+   41

O+   19

B-   28

A-   46

AB+   10

AB-   4

total   148


Probabilities:

a. A person has type A- blood: "p = Number(A-) \/ total = 46 \/ 148 = 0.31"

b. A person has type AB+ or type AB- blood: "p = (Number(AB+) + Number(AB-)) \/ total = (10 + 40) \/ 148 = 0.09"

c. A person has neither type B+ nor type A+ blood: "p = (total - Number(B+) - Number(A+)) \/ 148 = (148 - 0 - 41) \/ 148 = 0.72"